Solve the equation for exact solutions. 37) arcsin y- T6 3 A) B) {313+7 (31/3 - D) {6} 6
Answer #1

57 KM are sim (K) = x - k= sin(x) are sinly 7) = 7 - - Z Z = sin() y. B y + - 33**